### v1

@@ Abstract
Introduction – The Standard Models of particle physics and cosmology have been experimentally confirmed with high precision [1, 2]. Yet, open questions remain including the origin of the masses of neutrinos, the source of density fluctuations that evolved into the cosmic large-scale structure (LSS) and the fundamental natures of dark matter (DM) and dark energy (DE). Observations of the cosmic microwave background (CMB) radiation and the LSS are central to unravelling these mysteries.
@@
High-precision datasets such as CMB temperature T, polarization E,B and lensing $\phi$ anisotropies [3, 4, 5], baryon acoustic oscillations [BAO; 6] and supernovae magnitudes [SNe; 7] are consistent with the standard $\Lambda$ cold dark matter ($\Lambda$CDM) cosmological model. However, there is disagreement with some other datasets in the inference of cosmological parameters, which is either a sign of unmodeled systematics or deviation from $\Lambda$CDM. A local distance ladder using SNe and cepheid stars infers the expansion rate today $H_{0}$ to be $\sim 5\sigma$ higher than from an inverse distance ladder using CMB, BAO and SNe data in the $\Lambda$CDM model [8]. Further, some LSS data, in particular the clustering and lensing of photometric galaxy surveys like the Dark Energy Survey [DES; 9] and the Kilo-Degree Survey [KiDS; 10], show a $2-3\sigma$ preference, compared to CMB, BAO and SNe data, for a lower value of the cosmological parameter $S_{8}\equiv\sqrt{\frac{\Omega_{\mathrm{m}}}{0.3}}\sigma_{8}$, where $\Omega_{\mathrm{m}}$ is the total matter density and $\sigma_{8}$ is the amplitude of density fluctuations averaged over $8h^{-1}\,\mathrm{Mpc}$ scales today.
@@
CMB data are currently sensitive to the matter power spectrum for wavenumbers $k\lesssim 0.2h\,\mathrm{Mpc}^{-1}$ [Fig. 1; see also 4] through the generation of lensing anisotropies. A key probe of physics beyond $\Lambda$CDM is therefore the high-$k$ (small scale) matter power spectrum. Indeed, the $S_{8}$ discrepancy may be a hint of a small-scale deviation from the standard model as galaxy clustering and lensing data from DES and KiDS are modeled for $k\lesssim 1h\,\mathrm{Mpc}^{-1}$. Resolutions to the $S_{8}$ discrepancy center around scale-dependent changes to the growth of perturbations, e. g., ultra-light axion (ULA) DM [11], time-dependent changes, e. g., DM-DE interactions [12, 13], decaying DM [14, 15, 16, 17, 18], or stronger galactic feedback [19]. However, interpretation is complicated by uncertainties like non-linearities, the intrinsic alignment of galaxy shapes [20] and a lack of clarity as to the scales at which a discrepancy between CMB and LSS occurs (see below).
@@
The Lyman-alpha forest, quasar spectral absorption formed in the intergalactic medium (IGM), is a powerful probe of high-$k$, quasi-linear modes for $k\gtrsim 0.3h\,\mathrm{Mpc}^{-1}$ [Fig. 1; see also 21]. This probe therefore provides an independent test of consistency between CMB and LSS data uncomplicated by the problems outlined above, albeit with its own uncertainties regarding the astrophysical state of the IGM. This power arises because it is sourced by moderate IGM fluctuations (tracing underlying DM fluctuations with overdensities $\lesssim 10$) that are much less affected by non-linearities and galactic feedback than equivalent scales in galaxy surveys [e. g., 22].
@@
Data likelihoods – For a number of cosmological models ($\Lambda$CDM, massive neutrinos, power spectrum running, spatial curvature), all the cosmological information in the eBOSS flux power spectrum can be compressed to the amplitude $\left(\Delta_{\mathrm{lin}}^{2}\equiv\frac{k_{\mathrm{p}}^{3}P_{\mathrm{lin}}(k_{\mathrm{p}},z_{\mathrm{p}})}{2\pi^{2}}\right)$ and tilt $\left(n_{\mathrm{lin}}\equiv\frac{\mathrm{d}\mathrm{ln}P_{\mathrm{lin}}(k,z)}{\mathrm{d}\mathrm{ln}k}|_{k_{\mathrm{p}},z_{\mathrm{p}}}\right)$ of the linear matter power spectrum $P_{\mathrm{lin}}(k,z)$ evaluated at a central redshift $z_{\mathrm{p}}=3$ and wavenumber $k_{\mathrm{p}}=0.009\,\mathrm{s}\,\mathrm{km}^{-1}$. This compression holds for models that are sufficiently close to Einstein-de Sitter expansion for the redshifts probed by data and where the linear matter power spectrum is well approximated by a power-law relation for the wavenumbers probed by data [27, 28, 29, 30]. Further, when combining with external CMB data, it is accurate to use the marginal likelihood on $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}}$ [29]. In the Supplemental Material, we confirm that the compression holds for mixed ULA and WDM models. We therefore use a compressed 2D Gaussian likelihood derived from the eBOSS (Sloan Digital Sky Survey Data Release 14; SDSS DR14) Lyman-alpha forest flux power spectrum [24] and tabulated in Ref. [31]: the mean values of $\Delta_{\mathrm{lin}}^{2}$ and $n_{\mathrm{lin}}$ are respectively $0.310$ and $-2.340$, the standard deviations are respectively 0.020 and 0.006 and the correlation coefficient is 0.512 (Fig. 2). We give details in the Supplemental Material on the astrophysical model which is marginalized.
For CMB data, we use the Plik low-multipole $\ell$ TT and EE, $-\ell$ TTTEEE and conservative lensing $\phi\phi$ likelihoods from the Planck 2018 data release [1, 3]. We use BAO measurements from the CMASS and LOWZ galaxy samples of BOSS Data Release 12 at $z=0.38$, $0.51$ and $0.61$ [32], the 6dF Galaxy Survey (6dFGS) at $z=0.106$ [33] and SDSS Data Release 7 Main Galaxy Sample (SDSS DR7 MGS) at $z=0.15$ [34]; and the Pantheon+ catalog of uncalibrated SNe luminosity distances [7]. This compendium of CMB, BAO and SNe data constitutes a larger-scale complement to the smaller scales probed by the eBOSS flux power spectrum.
@@
Models and inference – We use a uniform prior distribution on baseline $\Lambda$CDM parameters: the Hubble constant $H_{0}$ [^1] For technical reasons, we sometimes instead use the angular size of the sound horizon $\theta_{\mathrm{s}}$., the physical baryon $\omega_{\mathrm{b}}$ and CDM $\omega_{\rm CDM}$ energy densities, the amplitude $A_{\mathrm{s}}$ and tilt $n_{\mathrm{s}}$ of the primordial power spectrum and the reionization optical depth $\tau$. Unless specified otherwise, we model free-streaming neutrinos as two massless and one massive neutrino with the sum of neutrino masses $\Sigma m_{\nu}=0.06$ eV. We consider four extensions of the $\Lambda$CDM model that suppress power on small scales to which the Lyman-alpha forest is sensitive with the following additional parameters (Table 1): (i) varying the sum of neutrino masses $\Sigma m_{\nu}$; (ii) the running of the primordial spectral index $\alpha_{\mathrm{s}}\equiv\frac{\mathrm{d}n_{\mathrm{s}}}{\mathrm{dln}k}$; (iii) ultra-light axions with particle mass $m_{\rm ULA}$ forming a fraction $f_{\mathrm{ULA}}$ of the DM; (iv) warm dark matter with particle mass $m_{\rm WDM}$ and temperature $T_{\rm WDM}$ (forming a fraction $f_{\mathrm{WDM}}$ of the DM). We use a uniform prior on $\Sigma m_{\nu},\alpha_{\mathrm{s}},\mathrm{log}[m_{\rm ULA}],f_{\rm ULA},m_{\rm WDM},T_{\rm WDM}$. We set prior ranges $\mathrm{log}[m_{\rm ULA}\,(\mathrm{eV})]\in[-27,-23],f_{\rm ULA}\in[0,0.05],m_{\rm WDM}\in[0,200]\,\mathrm{eV},\frac{T_{\rm WDM}}{T_{\rm CMB}}\in[0,1]$, where $T_{\mathrm{CMB}}$ is the CMB temperature, with the additional constraint $f_{\rm WDM}<0.05$. Restricting the DM fractions to no greater than 0.05 ensures that we satisy the condition for using the compressed likelihood while still optimizing the joint fit to CMB and Lyman-alpha forest data. We vary nuisance foreground parameters in the Planck likelihood and SNe calibrations in the Pantheon+ likelihood. The neutrino and running models were constrained by eBOSS Lyman-alpha forest data in Ref. [26] to which we compare our results below.
@@
WDM $1.63$27.08 $\mathrm{log}[m_{\mathrm{WDM}}(\mathrm{eV})]=1.35\pm^{0.60}_{0.30}$
$f_{\mathrm{WDM}}=0.0250\pm^{0.0021}_{0.0079}$
@@
We use the cosmological inference package MontePython-v3222https://github.com/brinckmann/montepython_public. [35, 36] interfaced with a modified version [^3] https://github.com/PoulinV/AxiCLASS. of the Einstein-Boltzmann solver CLASS [^4] https://lesgourg.github.io/class_public/class.html. [37, 38] in order to perform a Markov chain Monte Carlo (MCMC) sampling of the posterior distribution using a Metropolis Hasting algorithm. We model ULAs following the effective fluid approach from the modified Einstein-Bolztmann solver AxionCAMB [39] and implemented in AxiCLASS [^5] An improved fluid model was derived in Ref. [40] but the difference is negligible given the precision of our data. [41, 42]. We consider MCMC chains converged when the Gelman-Rubin statistic $|R-1|\lesssim 0.05$ [43]. To analyze MCMC chains and visualize posteriors, we use the GetDist package [44]. We find minimum $\chi^{2}$ values using the simulated annealing method described in the appendix of Ref. [45].
Results – Figure 2 shows the marginalized posterior of $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}},\sigma_{8},n_{\mathrm{s}}$. For the models that we consider, all the cosmological information in the eBOSS Lyman-alpha forest is contained in $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}}$ and so the posterior of those parameters is the same for each model. When the information in the flux power spectrum is projected to $\sigma_{8},n_{\mathrm{s}}$, which characterize the amplitude and tilt of the linear matter power spectrum at different redshifts and smaller wavenumbers, the relative constraint on these parameters is weaker than for $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}}$. As a consequence, there is broad agreement between CMB + BAO + SNe and Lyman-alpha forest data when considering only those parameters even under the $\Lambda$CDM model. In the Supplemental Material, we explore the use of prior information, in particular on $H_{0}$ in order to set the conversion from native velocity to spatial wavenumbers, but still find broad consistency with CMB + BAO + SNe. However, when projecting the information in Planck, BAO and SNe data to the larger pivot wavenumber that eBOSS probes, there is a $4.90\sigma$ tension [^6] We measure tension in the non-Gaussian marginalized $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}}$ constraints using the parameter difference posterior distribution as implemented in the tensiometer package [46]. under $\Lambda$CDM (Table 1) driven by a suppressed small-scale power spectrum amplitude and tilt. This tension is not alleviated when varying neutrino masses since the suppression is step-like at the wavenumbers to which eBOSS is sensitive (Fig. 1). However, the tension is reduced to $0.92\sigma$, $1.63\sigma$ and $0.56\sigma$ respectively for the one-parameter power spectrum running and two-parameter WDM and ULA models; these models suppress both power amplitude and spectral index.
@@
Fig. 3 shows joint posteriors of extended model parameters (full sets of 1D and 2D marginalized posteriors in the Supplemental Material). Increasing neutrino masses marginally suppresses the small-scale amplitude but cannot lower the small-scale tilt sufficiently. Increasing the deviation from no running can successfully lower the small-scale amplitude and tilt in agreement with Ref. [26] (see Supplemental Material for results where we simultaneously vary $\Sigma m_{\nu}$ and $\alpha_{\mathrm{s}}$). The ULA constraint, which disfavors the case with no axions, has a characteristic “w” shape. The WDM constraint shows an anti-correlation between particle mass and temperature with marginalized limits $\mathrm{log}[m_{\rm WDM}\,(\mathrm{eV})]=1.35\pm^{0.60}_{0.30}$ and $\frac{T_{\rm WDM}}{T_{\rm CMB}}=0.155\pm^{0.018}_{0.052}$, giving $f_{\rm WDM}=0.0250\pm^{0.0021}_{0.0079}$ (all mean and 68% c.l.). For both ULA and WDM, as the particle mass increases, the suppression wavenumber increases until it is larger than the eBOSS pivot and so doesn’t lower $\Delta_{\mathrm{lin}}^{2}$. For a given mass, there is a range of ULA fractions and WDM temperatures that moves the power spectrum transfer function to have the correct amplitude and tilt.
@@
We find that varying the masses of neutrinos cannot resolve this parameter discrepancy. This means that constraints on neutrino masses (when not varying any additional parameters) from Planck CMB and eBOSS Lyman-alpha forest are derived from datasets in significant tension and may not therefore be reliable; Ref. [26] presents limits when varying neutrino and running parameters (see also Supplemental Material). We set the first constraints on the ULA and WDM fractions from eBOSS Lyman-alpha forest; we find a preference for ULAs with $m_{\mathrm{ULA}}\sim 10^{-25}\,\mathrm{eV}$ and $f_{\mathrm{ULA}}\sim(1-5)\%$ improving the best fit by $\Delta\chi^{2}=-27.6$ over $\Lambda$CDM with no residual tension ($0.56\sigma$). Similarly, we find a contribution of WDM improves the fit by $\Delta\chi^{2}=-27.1$ with $m_{\rm WDM}\sim 90$ eV, $T_{\rm WDM}\simeq 0.19\,T_{\rm CMB}$ and reduced tension ($1.63\sigma$). With one fewer free parameter, running of the power spectrum reduces the tension to $0.92\sigma$ and improves the fit by $\Delta\chi^{2}=-25.6$ [see also 26].
@@
A key effect that changes the tilt of the matter power spectrum inferred from the Lyman-alpha forest is residual contamination from high-column-density (HCD) absorbers [25, 24, 26]. The eBOSS Collaboration found that using a more accurate model for residual HCDs [47, 48], calibrated on Illustris hydrodynamical simulations [49, 50], does not improve the flux power spectrum fit. Ref. [51] re-analyzed the eBOSS flux power spectrum with simulations including spatial fluctuations in helium reionization [52], finding a higher value of the power spectrum tilt that would disfavor running models relative to ULAs; although, their astrophysical model is not consistent between flux power spectrum and high-resolution spectrum IGM temperature [53] measurements. For upcoming Dark Energy Spectroscopic Instrument (DESI) data [54], where the statistical precision will reach percent level, it is crucial that the fidelity of IGM simulations further increases accounting for scale-dependent effects like HCDs. Coupled with higher-accuracy IGM simulations, it will be feasible to distinguish further between the models we have analyzed and astrophysical uncertainties like HCDs. LSS measurements of the matter power spectrum, e. g., the Rubin Observatory [55] and Simons Observatory [56], will be highly complementary.
@@
Models that reduce the eBOSS Lyman-alpha forest tension suppress power at higher wavenumbers $k\sim 10\,h\,\mathrm{Mpc}^{-1}$ that are probed by, e. g., higher-resolution Lyman-alpha forest [57, 58, 59, 60, 61, 62, 63] and Milky Way sub-structure [64, 65]. We anticipate that the $\sim 10\%$ suppression implied by eBOSS is consistent at the current sensitivity of higher-resolution flux power spectrum measurements from Keck and VLT (which have $\sim 25\%$ uncertainty); the XQ-100 Lyman-alpha forest dataset [60], which probes smaller scales than eBOSS, infers an even lower value of $n_{\mathrm{lin}}$, which may favor power spectrum running models. Hints of new physics were found in the HIRES/MIKE quasar sample [66]. We defer a detailed joint analysis of eBOSS and small-scale structure probes to future work.
@@
Acknowledgments – The authors thank Andreu Font-Ribera for valuable discussion. The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto. VP is supported by funding from the European Research Council (ERC) under the European Union’s HORIZON-ERC-2022 (grant agreement no. 101076865). VP is also supported by the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement no. 860881-HIDDeN. The authors acknowledge the use of computational resources from the Excellence Initiative of Aix-Marseille University (A*MIDEX) of the Investissements d’Avenir program and from the LUPM’s cloud computing infrastructure founded by Ocevu labex and France-Grilles.

### v3

@@ Abstract
Introduction – The Standard Models of particle physics and cosmology have been experimentally confirmed with high precision [1, 2]. Yet, open questions remain including the origin of the masses of neutrinos, the source of density fluctuations that evolved into the cosmic large-scale structure (LSS) and the fundamental natures of dark matter (DM) and dark energy (DE).
@@
High-precision datasets such as cosmic microwave background (CMB) temperature T, polarization E,B and lensing $\phi$ anisotropies [3, 4, 5], baryon acoustic oscillations [BAO; 6] and supernovae magnitudes [SNe; 7] are consistent with the standard $\Lambda$ cold dark matter ($\Lambda$CDM) cosmological model. However, there is disagreement with some other datasets in the inference of cosmological parameters, which is either a sign of unmodeled systematics or deviation from $\Lambda$CDM. A local distance ladder using SNe and cepheid stars infers the expansion rate today $H_{0}$ to be $\sim 5\sigma$ higher than from an inverse distance ladder using CMB, BAO and SNe data in the $\Lambda$CDM model [8]. Further, some LSS data, in particular the clustering and lensing of photometric galaxy surveys like the Dark Energy Survey [DES; 9] and the Kilo-Degree Survey [KiDS; 10], show a $2-3\sigma$ preference, compared to CMB, BAO and SNe data, for a lower value of the cosmological parameter $S_{8}\equiv\sqrt{\frac{\Omega_{\mathrm{m}}}{0.3}}\sigma_{8}$, where $\Omega_{\mathrm{m}}$ is the total matter density and $\sigma_{8}$ is the amplitude of density fluctuations averaged over $8h^{-1}\,\mathrm{Mpc}$ scales today.
@@
CMB data are currently sensitive to the matter power spectrum for wavenumbers $k\lesssim 0.2h\,\mathrm{Mpc}^{-1}$ [Fig. 1; see also 4] through the generation of lensing anisotropies. A key probe of physics beyond $\Lambda$CDM is therefore the high-$k$ (small scale) matter power spectrum. Indeed, the $S_{8}$ discrepancy may be a hint of a small-scale deviation from the standard model as galaxy clustering and lensing data from DES and KiDS are modeled for $k\lesssim 1h\,\mathrm{Mpc}^{-1}$. Resolutions to the $S_{8}$ discrepancy center around scale-dependent changes to the growth of perturbations, e. g., ultra-light axion (ULA) DM [11], time-dependent changes, e. g., DM-DE interactions [12, 13, 14], decaying DM [15, 16, 17, 18, 19], or stronger galactic feedback [20]. However, interpretation is complicated by uncertainties like non-linearities, the intrinsic alignment of galaxy shapes [21] and a lack of clarity as to the scales at which a discrepancy between CMB and LSS occurs (see below).
@@
The Lyman-alpha forest, quasar spectral absorption formed in the intergalactic medium (IGM), is a powerful probe of high-$k$, quasi-linear modes for $k\gtrsim 0.3h\,\mathrm{Mpc}^{-1}$ [Fig. 1; see also 22]. This probe therefore provides an independent test of consistency between CMB and LSS data uncomplicated by the problems outlined above, albeit with its own uncertainties regarding the astrophysical state of the IGM. This power arises because it is sourced by moderate IGM fluctuations (tracing underlying DM fluctuations with overdensities $\lesssim 10$) that are much less affected by non-linearities and galactic feedback than equivalent scales in galaxy surveys [e. g., 23].
@@
Data likelihoods – For a number of cosmological models ($\Lambda$CDM, massive neutrinos, power spectrum running, spatial curvature), all the cosmological information in the eBOSS flux power spectrum can be compressed to the amplitude $\left(\Delta_{\mathrm{lin}}^{2}\equiv\frac{k_{\mathrm{p}}^{3}P_{\mathrm{lin}}(k_{\mathrm{p}},z_{\mathrm{p}})}{2\pi^{2}}\right)$ and tilt $\left(n_{\mathrm{lin}}\equiv\frac{\mathrm{d}\mathrm{ln}P_{\mathrm{lin}}(k,z)}{\mathrm{d}\mathrm{ln}k}|_{k_{\mathrm{p}},z_{\mathrm{p}}}\right)$ of the linear matter power spectrum $P_{\mathrm{lin}}(k,z)$ evaluated at a central redshift $z_{\mathrm{p}}=3$ and wavenumber $k_{\mathrm{p}}=0.009\,\mathrm{s}\,\mathrm{km}^{-1}$. This compression holds for models that are sufficiently close to Einstein-de Sitter expansion for the redshifts probed by data and where the linear matter power spectrum is well approximated by a power-law relation for the wavenumbers probed by data [28, 29, 30, 31]. Further, when combining with external CMB data, it is accurate to use the marginal likelihood on $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}}$ [30]. In the Supplemental Material, we confirm that the compression holds for mixed ULA and WDM models. We therefore use a compressed 2D Gaussian likelihood derived from the eBOSS (Sloan Digital Sky Survey Data Release 14; SDSS DR14) Lyman-alpha forest flux power spectrum [25] and tabulated in Ref. [32] [^1] The likelihood of Ref. [25] is not publicly available and so Ref. [32] tabulated it (in their Table I) by fitting a 2D Gaussian to a $30\times 30$ grid of likelihood samples from Fig. 20 of Ref. [25]. There is a slight non-Gaussian feature that is not captured but, if the full likelihood were available, we do not anticipate that our conclusions would change. Further, Ref. [27] extends the analysis of Ref. [25] by testing extended cosmological models, but the data and likelihood are the same in both analyses; the astrophysical priors are different in the two analyses.: the mean values of $\Delta_{\mathrm{lin}}^{2}$ and $n_{\mathrm{lin}}$ are respectively $0.310$ and $-2.340$, the standard deviations are respectively 0.020 and 0.006 and the correlation coefficient is 0.512 (Fig. 2). We give details in the Supplemental Material on the astrophysical model which is marginalized. We also sometimes use a compendium of Planck 2018 CMB [1, 3], galaxy BAO [33, 34, 35] and Pantheon+ SNe [7] data that constitutes a larger-scale complement to the smaller scales probed by the eBOSS flux power spectrum; full details are given in the Supplemental Material.
@@
Models and inference – We consider four extensions to the baseline six-parameter $\Lambda$CDM model: (i) varying the sum of neutrino masses $\Sigma m_{\nu}$; (ii) the running of the primordial spectral index $\alpha_{\mathrm{s}}\equiv\frac{\mathrm{d}n_{\mathrm{s}}}{\mathrm{dln}k}$; (iii) ultra-light axions with particle mass $m_{\rm ULA}$ forming a fraction $f_{\mathrm{ULA}}$ of the DM; (iv) warm dark matter with particle mass $m_{\rm WDM}$ and temperature $T_{\rm WDM}$ (forming a fraction $f_{\mathrm{WDM}}$ of the DM). The linear matter power spectrum and other cosmological background observables are calculated for these models using the Einstein-Boltzmann solver AxiCLASS [36, 37]. We use a uniform prior on $\Sigma m_{\nu},\alpha_{\mathrm{s}},\mathrm{log}[m_{\rm ULA}],f_{\rm ULA},\mathrm{log}[m_{\rm WDM}],T_{\rm WDM}$ and the standard $\Lambda$CDM parameters. We restrict the ULA and WDM fractions to be no greater than 0.05 to ensure that we satisfy the condition for using the compressed likelihood while still optimizing the joint fit to CMB and Lyman-alpha forest data. More details are given on the models, prior and posterior sampling method in the Supplemental Material. The neutrino and running models are constrained by eBOSS Lyman-alpha forest data in Ref. [27] to which we compare our results below.
@@
WDM $1.34$27.08 $\mathrm{log}[m_{\mathrm{WDM}}(\mathrm{eV})]=1.01\pm^{0.30}_{0.44}$
$f_{\mathrm{WDM}}=0.0219\pm^{0.0030}_{0.0042}$
@@
Results – Figure 2 shows the marginalized posterior of $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}},\sigma_{8},n_{\mathrm{s}}$. For the models that we consider, all the cosmological information in the eBOSS Lyman-alpha forest is contained in $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}}$ and so the posterior of those parameters is the same for each model. When the information in the flux power spectrum is projected to $\sigma_{8},n_{\mathrm{s}}$, which characterize the amplitude and tilt of the linear matter power spectrum at different redshifts and smaller wavenumbers, the relative constraint on these parameters is weaker than for $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}}$. As a consequence, there is broad agreement between CMB + BAO + SNe and Lyman-alpha forest data when considering only those parameters even under the $\Lambda$CDM model. In the Supplemental Material, we explore the use of prior information, in particular on $H_{0}$ in order to set the conversion from native velocity to spatial wavenumbers, but still find broad consistency with CMB + BAO + SNe. However, when projecting the information in Planck, BAO and SNe data to the larger pivot wavenumber that eBOSS probes, there is a $4.90\sigma$ tension [^2] We measure tension in the non-Gaussian marginalized $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}}$ constraints using the parameter difference posterior distribution as implemented in the tensiometer package [38]. under $\Lambda$CDM (Table 1) driven by a suppressed small-scale power spectrum amplitude and tilt. This tension is not alleviated when varying neutrino masses since the suppression is step-like at the wavenumbers to which eBOSS is sensitive (Fig. 1). However, the tension is reduced to $0.92\sigma$, $1.34\sigma$ and $0.56\sigma$ respectively for the one-parameter power spectrum running and two-parameter WDM and ULA models; these models suppress both power amplitude and spectral index.
@@
Fig. 3 shows joint posteriors of extended model parameters (full sets of 1D and 2D marginalized posteriors in the Supplemental Material). Increasing neutrino masses marginally suppresses the small-scale amplitude but cannot lower the small-scale tilt sufficiently. Increasing the deviation from no running can successfully lower the small-scale amplitude and tilt in agreement with Ref. [27] (see Supplemental Material for results where we simultaneously vary $\Sigma m_{\nu}$ and $\alpha_{\mathrm{s}}$). The ULA constraint, which disfavors the case with no axions, has a characteristic “w” shape. The WDM constraint shows an anti-correlation between particle mass and temperature with marginalized limits $\mathrm{log}[m_{\rm WDM}\,(\mathrm{eV})]=1.01\pm^{0.30}_{0.44}$ and $\frac{T_{\rm WDM}}{T_{\rm CMB}}=0.194\pm^{0.043}_{0.074}$, where $T_{\mathrm{CMB}}$ is the CMB temperature, giving $f_{\rm WDM}=0.0219\pm^{0.0030}_{0.0042}$ (all mean and 68% c.l.). For both ULA and WDM, as the particle mass increases, the suppression wavenumber increases until it is larger than the eBOSS pivot and so doesn’t lower $\Delta_{\mathrm{lin}}^{2}$. For a given mass, there is a range of ULA fractions and WDM temperatures that moves the power spectrum transfer function to have the correct amplitude and tilt.
@@
We find that varying the masses of neutrinos cannot resolve this parameter discrepancy. This means that constraints on neutrino masses (when not varying any additional parameters) from Planck CMB and eBOSS Lyman-alpha forest are derived from datasets in significant tension and may not therefore be reliable; Ref. [27] presents limits when varying neutrino and running parameters (see also Supplemental Material) [^3] Indeed, this inconsistency applies for any model that does not induce suppression and running on small scales.. We set the first constraints on the ULA and WDM fractions from eBOSS Lyman-alpha forest; we find a preference for ULAs with $m_{\mathrm{ULA}}\sim 10^{-25}\,\mathrm{eV}$ and $f_{\mathrm{ULA}}\sim(1-5)\%$ improving the best fit by $\Delta\chi^{2}=-27.6$ over $\Lambda$CDM with no residual tension ($0.56\sigma$). Similarly, we find a contribution of WDM improves the fit by $\Delta\chi^{2}=-27.1$ with $m_{\rm WDM}\sim 10$ eV, $T_{\rm WDM}\simeq 0.19\,T_{\rm CMB}$ and reduced tension ($1.34\sigma$). With one fewer free parameter, running of the power spectrum reduces the tension to $0.92\sigma$ and improves the fit by $\Delta\chi^{2}=-25.6$ [see also 27].
A key effect that changes the tilt of the matter power spectrum inferred from the Lyman-alpha forest is residual contamination from high-column-density (HCD) absorbers [26, 25, 27]. The eBOSS Collaboration found that using a more accurate model for residual HCDs [39, 40], calibrated on Illustris hydrodynamical simulations [41, 42], does not improve the flux power spectrum fit. However, there are other systematic effects related to HCDs that are ignored in 1D Lyman-alpha forest analyses, but which our results suggest should now be investigated: e. g., the masking of HCDs in data is correlated to the large-scale structure; the efficiency of detection and masking is uncertain and so the residual population is unknown. Further, Ref. [43] demonstrated that mis-modeling of baryonic feedback from active galactic nuclei (AGN) can bias inference on $n_{\mathrm{s}}$ by $\sim 2\sigma$ using the Horizon-AGN simulations [44]. Given that the $S_{8}$ discrepancy can be explained by extremely strong feedback models [20], coupled with our new results, it is timely to consider the effect of feedback given different (more physical) hydrodynamical models, e. g., FIRE [45].
@@
Ref. [46] re-analyzed the eBOSS flux power spectrum with simulations including spatial fluctuations in helium reionization [47], finding a higher value of the power spectrum tilt that would disfavor running models relative to ULAs; although, their astrophysical model is not completely consistent between flux power spectrum and high-resolution spectrum IGM temperature [48] measurements [^4] This inconsistency may be relieved by varying cosmology when fitting the high-resolution data [Simeon Bird, private communication].. Ref. [49] also re-analyzed the data, finding a higher value of the tilt, albeit with a strong sensitivity in their inference to the prior distribution on the mean transmitted flux. Ref. [32] performed $\Lambda$CDM and early dark energy analyses using the compressed eBOSS likelihood; we use the compression as tabulated by Ref. [32] and therefore find our $\Lambda$CDM constraints are the same. For upcoming Dark Energy Spectroscopic Instrument (DESI) data [50], where the statistical precision will reach percent level, it is crucial that the fidelity of IGM simulations further increases accounting for scale-dependent effects like HCDs. Coupled with higher-accuracy IGM simulations, it will be feasible to distinguish further between the models we have analyzed and astrophysical uncertainties like HCDs. LSS measurements of the matter power spectrum, e. g., the Rubin Observatory [51] and Simons Observatory [52], will be highly complementary.
@@
Models that reduce the eBOSS Lyman-alpha forest tension suppress power at higher wavenumbers $k\sim 10\,h\,\mathrm{Mpc}^{-1}$ that are probed by, e. g., higher-resolution Lyman-alpha forest [53, 54, 55, 56, 57, 58, 59, 60] and Milky Way sub-structure [61, 62, 63]. We anticipate that the $\sim 10\%$ suppression implied by eBOSS is consistent at the current sensitivity of higher-resolution flux power spectrum measurements from Keck and VLT (which have $\sim 25\%$ uncertainty); the XQ-100 Lyman-alpha forest dataset [57], which probes smaller scales than eBOSS, infers an even lower value of $n_{\mathrm{lin}}$, which may favor power spectrum running models [^5] We do not use the compressed XQ-100 likelihood [32] as it has been compressed at the eBOSS pivot wavenumber despite XQ-100 data being sensitive to smaller scales. This approach necessarily means that a power-law matter power spectrum model has been assumed from eBOSS to XQ-100 scales, which does not apply to the extended cosmological models that we consider. We defer to future work a detailed analysis with dedicated simulations.. Hints of new physics were found in the HIRES/MIKE quasar sample [64, 65]. Ref. [66] identified a $\sim 3.3\sigma$ discrepancy in the inference of $\sigma_{8}$ between XQ-100 and HIRES/MIKE Lyman-alpha forest and galaxy cluster number counts from the South Pole Telescope Sunyaev-Zel’dovich effect [67]; they discuss HCD contamination in the context of this tension. We defer a detailed joint analysis of eBOSS and small-scale structure probes to future work.
@@
Acknowledgments – The authors thank Simeon Bird, Andreu Font-Ribera, Julien Lesgourgues, Riccardo Murgia and Nils Schöneberg for valuable comments. The authors also thank Cosmology Talks mini-workshops where elements of this project were conceived. KKR is supported by an Ernest Rutherford Fellowship from the UKRI Science and Technology Facilities Council (grant number ST/Z510191/1). The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto. VP is supported by funding from the European Research Council (ERC) under the European Union’s HORIZON-ERC-2022 (grant agreement no. 101076865). VP is also supported by the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement no. 860881-HIDDeN. The authors acknowledge the use of computational resources from the Excellence Initiative of Aix-Marseille University (A*MIDEX) of the Investissements d’Avenir program and from the LUPM’s cloud computing infrastructure founded by Ocevu labex and France-Grilles.